# Who Made This? Fake Detection and Source Attribution with Diffusion Features

### v1

@@ Abstract
The rapid progress of generative diffusion models has enabled the creation of synthetic images that are increasingly difficult to distinguish from real ones, raising concerns about authenticity, copyright, and misinformation. Existing supervised detectors often struggle to generalize across unseen generators, requiring extensive labeled data and frequent retraining. We introduce FRIDA (Fake-image Recognition and source Identification via Diffusion-features Analysis), a lightweight framework that leverages internal activations from a pre-trained diffusion model for deepfake detection and source generator attribution. A k-nearest-neighbor classifier applied to diffusion features achieves state-of-the-art cross-generator performance without fine-tuning, while a compact neural model enables accurate source attribution. These results show that diffusion representations inherently encode generator-specific patterns, providing a simple and interpretable foundation for synthetic image forensics.
Index Terms —
Overview of our framework for deepfake classification and attribution. Given an input image, our Image Prototyping module extracts image prototype from the U-Net of a Stable Diffusion Model. This compact representation is then used for two downstream tasks: (i) Deepfake classification via a k-NN classifier to determine whether the image is real or fake, and (ii) Deepfake attribution via an MLP to identify the generative model (e.g., GLIDE, Stable Diffusion, BigGAN) responsible for the fake.
@@ 1 Introduction
The outstanding capabilities of recent generative models to synthesize media with exceptional fidelity and variety [1, 16, 4, 36] have accelerated the proliferation of AI‑generated content across domains, including medical imaging [23, 47, 42], education [53, 51], marketing [20, 12], and robotics [50]. This rapid progress has been largely driven by diffusion-based generators — including Stable Diffusion Model (SDM) [36], DALL-E  [35], and Imagen [37] — which produce highly realistic and semantically coherent visuals from simple text prompts. This increasing accessibility of advanced image generation tools through user-friendly interfaces poses considerable risks, including the creation of biased content and the violation of intellectual property rights. To mitigate these risks, robust detection and attribution methods are required to reliably identify synthetic media and determine their source models [39, 13, 29].
Early fake detection methods relied on supervised learning pipelines [52, 59, 29], which depend on vast labelled datasets and intensive computation. Approaches for source generator attribution pushed these requirements even further, as they must learn generator-specific fingerprints [54, 8, 46]. However, as generative models evolve rapidly, the continual need for data collection and retraining makes classifier-based detectors hard to scale, underscoring the need for lightweight, data-efficient methods that generalize across both known and emerging generators.
@@
An opportunity to move beyond task-specific classifiers has emerged with the rise of large-scale foundation models based on Vision Transformers (ViTs) [19], e.g., CLIP [34] and DINO [9]. Pre-trained on massive datasets, these models learn general, transferable representations that have shown remarkable effectiveness in downstream tasks [24, 45] — including deepfake detection [31, 40, 44] — when their extracted features are used for classification.
@@
Similarly, diffusion-based generative models, though designed for image generation, have proven to be remarkably effective feature encoders. Recent studies have shown that pre-trained SDMs provide rich semantic information about image content [5, 6], enabling strong performance in applications such as semantic segmentation [3, 15], and other vision-related problems — substantially reducing the computational cost of training large models from scratch. At the same time, diffusion models have been increasingly used for fake detection. Some methods identify diffusion-generated images through reconstruction error [49], while others analyse latent trajectories via diffusion inversion to enhance generalisation across unseen generators [10, 43]. Based on the principle that images from different sources exhibit distinct distributions, Zhong et al. [57] employ a SDM as a denoising tool to identify fake images by detecting the resulting artifacts in their reconstructed versions. However, these approaches typically require repeated diffusion or inversion steps, making them computationally expensive and less practical for large-scale or real-time detection scenarios.
@@
Inspired by these recent advances, we explore whether the internal features of a pre-trained diffusion-based generative model can effectively separate real from synthetic images and, beyond detection, enable attribution to the source generator. To this end, we introduce a lightweight and data-efficient framework that operates entirely within the latent feature space of SDMs at the last inference step. Our approach detects deepfakes by applying a training-free $k$-Nearest Neighbours ($k$-NN) classifier to latent representations extracted from the SDM. This approach achieves state-of-the-art performance on the GenImage benchmark  [59] for fake image detection. This strategy eliminates the need for costly inversion or fine-tuning and shows strong generalisation to unseen generators. We further extend our method to source-model attribution, showing that a lightweight Multi-Layer Perceptron (MLP) trained on the SDM latent features can accurately identify the generator responsible for a synthetic image.
@@
    We show that latent features extracted from specific layers of a pre-trained SDM are highly discriminative for both fake image detection and source model attribution.
@@
    We benchmark our fake image detection approach on data generated by eight different models, achieving state-of-the-art performance with a novel, lightweight $k$-NN-based framework that is entirely training-free. Our method requires only a small support set, generalises effectively to unseen generators, and adapts to new data without any fine-tuning.
@@
    A lightweight neural classifier trained on the SDM latent representation proves to be highly effective in source model attribution, suggesting the presence of generator-specific characteristics in the features. We use SHAP (SHapley Additive exPlanations) [28] to further investigate the capability of the SDM to encode these model-specific signatures.
The remainder of this paper is organized as follows:  Section 2 reviews relevant literature on synthetic image detection and attribution.  Section 3 describes the GenImage dataset and the procedure used to extract the latent representation of each image using the SDM.  Section 4 describes the experimental setup, while  Section 5 presents and discusses the obtained results. Finally,  Section 6 concludes the paper with a summary of our findings and outlines future research directions.
@@ 2 Related Work
Early research on detecting GAN-generated content showed that synthetic images contain characteristic frequency patterns and texture artifacts. Zhang et al. [56] introduced AutoGAN, a frequency-domain simulator of upsampling artifacts, and trained a spectrum-based classifier to detect periodic Fourier patterns without access to the generator. Similarly, Wang et al. [46] showed that CNN-generated images share common low-level artifacts across architectures and datasets. Their ResNet-based detector, trained on ProGAN, generalized remarkably well to unseen generators such as StyleGAN and BigGAN, revealing universal CNN fingerprints that persist across models. Gram-Net [26] models global texture correlations via multi-layer Gram matrices, improving robustness to compression and noise and enabling better cross-GAN generalization by leveraging long-range texture cues. Qian et al. [33] advanced this line of work with F3-Net, a two-branch frequency-aware CNN classifier combining frequency decomposition and local statistics via cross-attention, outperforming spatial-domain detectors on a deepfake faces dataset even under strong compression.
@@
Following the success of diffusion-based models, several works have focused on detecting diffusion-generated images. DE-FAKE [38] is a hybrid two-branch architecture that exploits image-text consistency between captions and visual content to detect synthetic images. DIRE [49] detects diffusion-generated images by inverting an input into the diffusion latent space and then reconstructing it through the full reverse denoising trajectory; the discrepancy between the original and reconstructed signals is used for detection. Similarly, LaRE2 [29] estimates a Latent Reconstruction Error using only a single-step reconstruction in latent space, using a module that aligns and refines features across spatial and channel dimensions. A related line of work [48] also leverages a pre-trained diffusion model, extracting multi-timestep responses under the hypothesis that synthetic images, which lie outside the natural image manifold, exhibit distinctive denoising behavior. ESIDE [52] incorporates frequency perturbations into diffusion inversion, training an ensemble of CLIP-based classifiers on noised representations to improve robustness and interpretability. More recently, LATTE [43] introduces a Latent Trajectory Embedding framework that explicitly models the temporal evolution of latent representations across denoising steps. By aggregating multi-timestep latent features through joint visual–latent refinement, LATTE captures dynamic generation cues beyond reconstruction error, achieving strong cross-generator and cross-dataset generalization. An alternative approach [57] treats detection as an anomaly detection task. They learn the low-level feature distribution of real images by training an extractor to spot pixel-level differences between original images and their denoised counterparts, effectively identifying generated content that falls outside this learned distribution.
@@
Recent research explores semantic alignment and open-set detection to generalize across diverse generators. Zhu et al. [58] propose MAID, a framework-agnostic attribution method that extracts Diffusion Model Activations (DMA) by treating pre-trained diffusion models as denoising autoencoders. These activations encode model-specific patterns without requiring white-box access or prompts, supporting both detection and attribution. SemGIR [55] employs semantic-guided image regeneration: a candidate image is captioned, regenerated via text-to-image synthesis, and compared with its reconstruction using CLIP-based encoders. This forces the detector to focus on generator-specific artifacts rather than prompt semantics and achieves strong cross-generator generalization. Cioni et al. [14] move beyond frequency-based fingerprints by leveraging intermediate representations of large ViT-based models such as CLIP and DINOv2. Their open-set attribution framework combines linear probing and k-nearest neighbors with confidence- or distance-based rejection to identify images from unseen generators, providing a unified and retraining-free solution across GAN and diffusion sources.
3 Materials and Methods
3.1 GenImage Dataset
@@
GenImage [59] is a large-scale benchmark designed to detect generated imagery, composed of both synthetic and real images sourced from ImageNet [17]. The dataset contains roughly 1.35 million synthetic images produced by eight distinct generative models: BigGAN [7], GLIDE [32], VQDM [21], SDM (v1.4 and v1.5)[36], ADM[18], Midjourney [1], and Wukong [2]. Both the real and synthetic images cover the 1,000 ImageNet classes, with the synthetic portion providing approximately 1,350 images per class (1,300 for training and 50 for testing). The generated images were created using simple text prompts following the template “photo of [class]” and have resolutions ranging from $128\times 128$ to $1024\times 1024$, depending on the source model. In this study, we utilize a subset of the training images. For each generator, we randomly sampled 10,000 real and 10,000 synthetic images (10 images per ImageNet class). These subsets are then partitioned into training ($80\%$) and validation ($20\%$) sets. In the rest of the paper, we will refer to these two sets as “training subset” and “validation set”. Finally, the original GenImage test set was used as a held-out evaluation set.
@@ 3.2 Image Prototype Extraction
We exploit a pre-trained SDM v1.5111https://github.com/hkproj/pytorch-stable-diffusion and we use the following procedure for extracting internal features prototype from a specific layer (see Figure 1).
@@ 4 Experimental Setup
This section details the experimental setup used in our study. First, we compare features from different layers of the diffusion U-Net to identify which layer is most effective at distinguishing real from synthetic content. We then describe the experimental setups for our two primary tasks: fake image detection (Section 4.2) and source generator attribution (Section 4.3).
@@ 4.1 U-Net Layer Selection by Linear Probing
To identify which layers provide the most effective features for deepfake detection, we extract latent representations from the encoder, bottleneck, and decoder layers of the U-Net at multiple spatial resolutions ($64\times 64$, $32\times 32$, $16\times 16$, and $8\times 8$). Following the GenImage protocol [59], we train eight classifiers, one for each generator in the dataset, and evaluate each classifier on images from all other generators. In particular, we adopt a linear probing setup — commonly used to evaluate the quality and separability of learned representations in self-supervised and foundation models  [11, 22, 34] — by training a linear classifier with sigmoid activation on the latent prototypes from each layer. Each classifier is trained on the training subset (see Section 3.1) with real and synthetic images from a given generator. All models are trained with the AdamW optimizer [27] using a learning rate of $3\times 10^{-4}$. Training is early stopped based on validation accuracy, with a patience of 15 epochs. To assess generalization, we adopt a cross-generator evaluation: following common practice [59], each classifier trained on a specific generator is evaluated on the fake detection task using its own validation set (seen data distribution) as well as the validation sets of the seven remaining generators (unseen data distributions). The U-Net layer whose features yield the highest average cross-generator accuracy is then selected and used for all subsequent experiments.

### v2

@@ Abstract
The rapid rise of generative models has yielded synthetic images of striking realism, blurring the line between real and fake content. As novel models proliferate, detectors must go beyond mere fake identification to robustly generalise across unseen generators and synthetic content. We introduce FRIDA (Fake image Recognition and source Identification via Diffusion features Analysis), a lightweight, data-efficient framework that uses features from a pre-trained Stable Diffusion Model to detect and attribute AI-generated images. Through an in-depth analysis of how data from different generators are encoded across diffusion U-Net layers, we propose a method that (i) detects synthetic images using a training-free $k$-Nearest Neighbour approach and (ii) performs source model attribution via a compact neural classifier. On the GenImage benchmark, FRIDA achieves state-of-the-art cross-generator detection with limited data while maintaining robust source model attribution capabilities. These results establish diffusion features as a reliable framework for AI-generated image forensics.
Overview of our framework for fake image classification and source model attribution. Given an input image, our Image Prototyping module extracts image prototype from the U-Net of a Stable Diffusion Model. This compact representation is then used for two downstream tasks: (i) fake image classification via a $k$-NN classifier to determine whether the image is real or fake, and (ii) source model attribution via an MLP to identify the generative model (e.g., GLIDE, Stable Diffusion, BigGAN) responsible for the fake.
@@ 1 Introduction
The outstanding capabilities of recent generative models to synthesize media with exceptional fidelity and variety [midjourney, nanobanana, betker2023improving, Rombach_2022_CVPR] have accelerated the proliferation of AI‑generated content across domains, including medical imaging [KAZEROUNI2023102846, WANG2025102593, 10937272], education [xia2024scoping, wu2025systematic], marketing [grewal2025generative, cillo2025generative], and robotics [wolf2025diffusion]. This rapid progress has been largely driven by diffusion-based generators—including Stable Diffusion Model (SDM) [Rombach_2022_CVPR], DALL-E  [ramesh2021zero], and Imagen [saharia2022photorealistic]—which produce highly realistic and semantically coherent visuals from simple text prompts. This increasing accessibility of advanced image generation tools through user-friendly interfaces poses considerable risks, including the creation of biased content and the violation of intellectual property rights. To mitigate these risks, robust detection and attribution methods are required to reliably identify synthetic media and determine their source models [sha2023fake, luo2024lare, cioni2024clip]. Early fake detection methods relied on supervised learning pipelines [wu2025explainable, zhu2023genimage, luo2024lare], which depend on vast labelled datasets and intensive computation. Approaches for source model attribution pushed these requirements even further, as they must learn generator-specific fingerprints [yu2019attributing, bui2022repmix, wang2020cnn]. However, as generative models evolve rapidly, the continual need for data collection and retraining makes classifier-based detectors hard to scale, underscoring the need for lightweight, data-efficient methods that generalize across both known and emerging generators.
@@
An opportunity to move beyond task-specific classifiers has emerged with the rise of large-scale foundation models based on Vision Transformers (ViTs) [dosovitskiy2021an], e.g., CLIP [radford2021learning] and DINO [caron2021emerging]. Pre-trained on massive datasets, these models learn general, transferable representations that have shown remarkable effectiveness in downstream tasks [kerssies2025your, wang2024enhancing]—including deepfake detection [nguyen2024exploring, smeu2025declip, wang2025fsfm]—when their extracted features are used for classification.
@@
Similarly, diffusion-based generative models, though designed for image generation, have proven to be remarkably effective feature encoders. Recent studies have shown that pre-trained SDMs provide rich semantic information about image content [BONECHI20245244, BONECHI2025128846], enabling strong performance in applications such as semantic segmentation [baranchuk2021label, corradini2024freeseg], and other vision-related problems—substantially reducing the computational cost of training large models from scratch. At the same time, diffusion models have been increasingly used for fake detection. Some methods identify diffusion-generated images through reconstruction error [wang2023dire], while others analyse latent trajectories via diffusion inversion to enhance generalisation across unseen generators [cazenavette2024fakeinversion, vasilcoiu2025latte]. Based on the principle that images from different sources exhibit distinct distributions, Zhong et al. [zhong2025beyond] employ a SDM as a denoising tool to identify fake images by detecting the resulting artifacts in their reconstructed versions. However, these approaches typically require repeated diffusion or inversion steps, making them computationally expensive and less practical for large-scale or real-time detection scenarios.
@@
Inspired by these recent advances, we explore whether the internal features of a pre-trained diffusion-based generative model can effectively separate real from synthetic images drawn by various generators and, beyond detection, enable attribution to the source generator. To this end, we propose FRIDA, a data-efficient framework for detecting fake images. The core of our method is to repurpose a pre-trained SDM (v1.5) at the last inference step as a feature extractor. Then, these features are employed by a simple $k$-Nearest Neighbours ($k$-NN) classifier, showing state-of-the-art results on the GenImage benchmark [zhu2023genimage]. This strategy eliminates the need for costly inversion or fine-tuning and shows strong generalisation to unseen generators. We further extend our method to source model attribution, showing that a lightweight Multi-Layer Perceptron (MLP) trained on the SDM latent features can accurately identify the generator responsible for a synthetic image.
@@
    We show that latent features extracted from specific layers of a pre-trained SDM are highly discriminative for fake image detection. Furthermore, we benchmark the efficacy of SDM as a feature extractor against prominent ViT-based backbones (CLIP and DINO), establishing that SDM-based representations achieve superior detection performance.
@@
    We evaluate our fake image detection approach on data generated by eight different models, achieving state-of-the-art performance on GenImage dataset with a novel $k$-NN-based framework that is entirely training-free. We further demonstrate the robustness of our method by testing on a challenging unseen subset comprising state-of-the-art generators, including Flux and SDv3.5. Our method requires only a small support set, generalises effectively to unseen generators, and adapts to new data without any fine-tuning.
@@
    We evaluate the robustness of FRIDA against varying levels of image perturbations, including Gaussian noise and JPEG compression, to simulate real-world distortions.
  4. 4.
    A lightweight neural classifier trained on the SDM latent representation proves to be highly effective in source model attribution, suggesting the presence of generator-specific characteristics in the features. We use SHAP (SHapley Additive exPlanations) [lundberg2017unified] to further investigate the capability of the SDM to encode these model-specific signatures.
The remainder of this paper is organized as follows:  Section˜2 reviews relevant literature on synthetic image detection and attribution. Section˜3 presents the benchmark dataset (GenImage) alongside an auxiliary dataset of recent generators, and outlines the methodology for extracting per-image latent representations using the SDM.  Section˜4 describes the experimental setup, while  Section˜5 presents and discusses the obtained results. Finally,  Section˜6 concludes the paper with a summary of our findings and outlines future research directions.
@@ 2 Related Work
Early research on detecting GAN-generated content showed that synthetic images contain characteristic frequency patterns and texture artifacts. Zhang et al. [zhang2019detecting] introduced AutoGAN, a frequency-domain simulator of upsampling artifacts, and trained a spectrum-based classifier to detect periodic Fourier patterns without access to the generator. Similarly, Wang et al. [wang2020cnn] showed that CNN-generated images share common low-level artifacts across architectures and datasets. Their ResNet-based detector, trained on images from ProGAN, generalized remarkably well to unseen generators such as StyleGAN and BigGAN, revealing universal CNN fingerprints that persist across models. Gram-Net [liu2020global] models global texture correlations via multi-layer Gram matrices, improving robustness to compression and noise and enabling better cross-GAN generalization by leveraging long-range texture cues. Qian et al. [qian2020thinking] advanced this line of work with F3-Net, a two-branch frequency-aware CNN classifier combining frequency decomposition and local statistics via cross-attention, outperforming spatial-domain detectors on a deepfake faces dataset even under strong compression.
@@
Following the success of diffusion-based models, several works have focused on detecting images generated by such models. DE-FAKE [defake2023] is a hybrid two-branch architecture that exploits image-text consistency between captions and visual content to detect synthetic images. DIRE [wang2023dire] detects diffusion-generated images by inverting an input into the diffusion latent space and then reconstructing it through the full reverse denoising trajectory; the discrepancy between the original and reconstructed signals is used for detection. Similarly, LaRE2 [luo2024lare] estimates a Latent Reconstruction Error using only a single-step reconstruction in latent space, using a module that aligns and refines features across spatial and channel dimensions. A related line of work [diffimplicit2025] also leverages a pre-trained diffusion model, extracting multi-timestep responses under the hypothesis that synthetic images, which lie outside the natural image manifold, exhibit distinctive denoising behavior. ESIDE [wu2025explainable] incorporates frequency perturbations into diffusion inversion, training an ensemble of CLIP-based classifiers on noised representations to improve robustness and interpretability. More recently, LATTE [vasilcoiu2025latte] introduces a Latent Trajectory Embedding framework that explicitly models the temporal evolution of latent representations across denoising steps. By aggregating multi-timestep latent features through joint visual-latent refinement, LATTE captures dynamic generation cues beyond reconstruction error, achieving strong cross-generator and cross-dataset generalization. An alternative approach [zhong2025beyond] treats detection as an anomaly detection task. It learns the low-level feature distribution of real images by training an extractor to spot pixel-level differences between original images and their denoised counterparts, effectively identifying generated content that falls outside this learned distribution.
@@
To generalize across diverse generators, recent research explores semantic alignment, which disentangles generator-specific artifacts from image semantics, and open-set detection, which extends attribution to generators unseen at training time. Zhu et al. [zhu2025maid] propose MAID, a framework-agnostic attribution method that extracts diffusion model activations by treating pre-trained diffusion models as denoising autoencoders. These activations encode model-specific patterns without requiring white-box access or prompts, supporting both detection and attribution. SemGIR [semgir2024] employs semantic-guided image regeneration: a candidate image is captioned, regenerated via text-to-image synthesis, and compared with its reconstruction using CLIP-based encoders. This forces the detector to focus on generator-specific artifacts rather than prompt semantics and achieves strong cross-generator generalization. Cioni et al. [clipfeat2024] move beyond frequency-based fingerprints by leveraging intermediate representations of large ViT-based models such as CLIP and DINO. Their open-set attribution framework combines linear probing and $k$-nearest neighbors with confidence-based or distance-based rejection to identify images from unseen generators, providing a unified and retraining-free solution across GAN and diffusion sources.
3 Dataset Overview and Image Prototyping
3.1 Datasets
GenImage Dataset.
@@ 3.1 Datasets
GenImage [zhu2023genimage] is a large-scale benchmark designed to detect generated imagery, composed of both synthetic and real images sourced from ImageNet [deng2009imagenet]. The dataset contains roughly 1.35 million synthetic images produced by eight distinct generative models: BigGAN [brock2018large], GLIDE [nichol2022glide], VQDM [gu2022vector], SDM (v1.4 and v1.5) [Rombach_2022_CVPR], ADM [dhariwal2021diffusion], Midjourney [midjourney], and Wukong [wukong]. Both the real and synthetic images cover the 1,000 ImageNet classes, with the synthetic portion providing 1,350 images per class (1,300 for training and 50 for testing). The generated images were created using simple text prompts following the template “photo of [class]” and have resolutions ranging from $128\times 128$ to $1024\times 1024$, depending on the source model. In this study, we utilize a subset of the training images. For each generator, we randomly sampled 10,000 real and 10,000 synthetic images (10 images per ImageNet class). These subsets are then partitioned into training ($80\%$) and validation ($20\%$) sets. In the rest of the paper, we will refer to these two sets as “training subset” and “validation set”. Finally, the original GenImage test set was used as a held-out evaluation set.
Other Datasets.
To evaluate generalization beyond GenImage, we assembled an alternative, out-of-distribution (OOD) test set by extracting 500 synthetic images per generator and 500 real images per dataset from different sources. In particular, fake images are sampled from the Qwen-Image-Self-Generated-Dataset [wu2025qwenimagetechnicalreport] and from five generators of the OpenFake dataset [livernoche2025openfake] (i.e., Flux 1.1 Pro, Flux 1.0 Dev, GPT Image 1, Ideogram 3.0, and SDv3.5). The real images are sampled from LAION 400M [schuhmann2021laion] and Pascal VOC [Everingham15].
3.2 The GenImage Evaluation Protocol
The cross-generator evaluation protocol used in this work follows the one in  [zhu2023genimage]. To assess generalization, a binary real-vs-fake classifier is trained on one generator-specific subset and then tested—without any adaptation—on evaluation data from both the same generator (seen data distribution) and from the other seven generators (unseen data distributions).
@@ 3.3 Image Prototype Extraction
We adopt a pre-trained SDMv1.5111https://github.com/hkproj/pytorch-stable-diffusion and we use the following procedure for extracting internal features prototype from a specific layer (see Figure˜1).
@@ 4 Experimental Setup
This section details the experimental setup used in our study. First, we compare features from different layers of the diffusion U-Net to identify which layer is most effective at distinguishing real from synthetic content (Section˜4.1). We then describe the experimental setups for our two primary tasks: fake image detection (Section˜4.2) and source model attribution (Section˜4.3). Finally, in Section˜4.4, we present additional studies using different backbones, out-of-distribution datasets and various image perturbations.
@@ 4.1 U-Net Layer Selection by Linear Probing
To identify which layers provide the most effective features for fake image detection, we extract latent representations from the encoder, bottleneck, and decoder layers of the U-Net at multiple spatial resolutions ($64\times 64$, $32\times 32$, $16\times 16$, and $8\times 8$). We adopt a linear probing setup—commonly used to evaluate the quality and separability of learned representations in self-supervised and foundation models  [chen2020simple, he2020momentum, radford2021learning]—by training a linear classifier with sigmoid activation on the latent prototypes from each layer. Each classifier is trained on the training subset (see Section˜3.1) with real and synthetic images from a given generator. All models are trained with the AdamW optimizer [loshchilov2018decoupled] using a learning rate of $3\times 10^{-4}$. Training is early stopped based on validation accuracy, with a patience of 15 epochs. To assess generalization, we perform the cross-generator evaluation described in Section˜3.2 using the eight validation sets. The U-Net layer whose features yield the highest average cross-generator validation accuracy is then selected and used for all subsequent experiments.